How do you solve these problems
2. How do you solve 4x +7y=14
x=41-8y

Answer by richwmiller(17219)

You are allowed one problem at a time.
substitute 41-8y for x in the first equation
4*(41-8y)+7y=14
y=6
x=41-8*(6)
x=41-48
x=-7
BTW What makes you think this is a travel word problem?
The category may not seem important to you but it is to the website.